OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x))

OpenStudy (anonymous):

f(3x+2)=4(3x+2)+1 =12x+8+1 =12x+9 or =3(4x+3)

OpenStudy (anonymous):

thanks man